OpenStudy (anonymous):

I DONT EVEN KNOW WHERE TO START :( Give an example of a rational function that has a horizontal asymptote at y = 0 and a vertical asymptote at, x = 2 and x = 1.

OpenStudy (anonymous):

hint, the vertical asymptotes will occur where the denominator is zero

OpenStudy (anonymous):

so make a quadratic with zeros at \(1\) and \(2\) is it clear how to do that ?

OpenStudy (anonymous):

well what do you mean by quadratic for the zeros

OpenStudy (anonymous):

what polynomial has zeros at \(x=1\) and \(x=2\) ? hint, write it in factored form

OpenStudy (anonymous):

factored for as in x-1=0 and x-2=0??

OpenStudy (anonymous):

right but actually i meant as \((x-1)(x-2)\) that is your denominator

OpenStudy (anonymous):

so (x-1)(x-2) is all my denominator, but i dont get how i would come up with the numerator

OpenStudy (anonymous):

would it just be one?

OpenStudy (anonymous):

what is the degree of your denominator?

OpenStudy (anonymous):

yeah you could use \(1\) that would work \[\frac{1}{(x-1)(x-2)}\] would work

OpenStudy (anonymous):

so would any number up top

OpenStudy (anonymous):

and since your denominator is of degree 2, a degree1 polynomial up top would work as well for example \[\frac{x}{(x-1)(x-2)}\]

OpenStudy (anonymous):

ohh because of the rule where if the numerator degree is < the denom degree the asymptote is 0

OpenStudy (anonymous):

so this would count as a rational equation for the question. im not missing anything right? 1/(x-1)(x-2)